A whole chicken sells for $7.06 and has a mass of 1.5 kg. A beef shank sells for $10.00 with a mass of 2.5 kg. Compare the per pound cost for each item. (1 kg= 2.2 lb)
Found 2 solutions by RAY100, stanbon:
Answer by RAY100(1637)

Chicken,,,$ 7.06 / 1.5 kg * (1kg/2.2#) = $ 2.14 / #
.
Beef,,,,,,$ 10.00 / 2.5 kg * (1kg/2.2#) = $ 1.82 /#
.

Answer by stanbon(75887)

A whole chicken sells for $7.06 and has a mass of 1.5 kg. A beef shank sells for $10.00 with a mass of 2.5 kg. Compare the per pound cost for each item.
(1 kg= 2.2 lb)
1.5 kg = 1.5*2.2 lb = 3.3 lb
-
2.5 kg = 2.5*2.2 lb = 5.5 lb
-----------------------------------------
per lb price for chicken = 7.06/3.3 lb = $2.139 per lb.
-----------------------------------------
per lb price for beef = 10.00/5.5 lb = $1.82 per lb.
==============================================
Cheers,
Stan H.